# A unique Neutrino Mass Matrix Texture under Exponential Parametrization

### v2

@@ 4.2 Important Relations
$\frac{v_{\chi}^{2}}{v_{\eta}}$ $=$ $\frac{\Lambda^{2}y_{c}\left(b_{1}+\sqrt{3}b_{2}-g\right)^{2}}{6v^{2}y_{2}^{2}(2b_{1}-2ib_{2}+f)},$ (5)
$v_{\xi}$ $=$ $-\frac{6\sqrt{3}v^{2}y_{1}^{2}}{\sqrt{3}ay_{a}+2\sqrt{3}b_{1}y_{a}+6b_{2}y_{a}},$ (6)
where, $b_{1}=r\cos\theta$ and $b_{2}=r\sin\theta$. We expect that these relations, somehow, will help to understand the numerical domain of the model parameters more closely. The parameter space of the elements, $v_{1}/v_{2}$, $|v_{\chi}^{2}/v_{\eta}|$ and $|v_{\xi}|$ are presented in Fig. 5). For this, we allow the cut-off scale and the Yukawa couplings to vary randomly in the bounds: $(10^{3}-10^{16})$ GeV and $(0.01-0.99)$ respectively. Needless to mention, the $v_{H}$ is taken as $246$ GeV.
Figure 5: Shows the correlation plots between (a) $v_{1}/v_{2}$ vs $|v_{\xi}|$. (b) $log(|v_{\xi}|)$ vs $log(|v_{\chi}^{2}/v_{\eta}|)$. (c) $log(\Lambda)$ vs $log(|v_{\xi}|)$. The red points signify texture prediction (TP) while the blue data points signify the corrections from model sector(MP).
@@ 4.3 The Scalar Sector
There are seven complex scalar fields in the model and, the associated $SU(2)_{L}\times U(1)_{Y}\times A_{4}\times Z_{10}$ invariant potential, is shown explicitly in B. We see that there are seven $\mu^{2}$’s and fifty one $\lambda$’s in the potential expression. Thus it enhances the liberty to choose the desired vacuum alignments for the scalar fields. The eleven minimisation conditions in support of the vacuum alignments as discussed in section 4.1 are elaborated in B.
From the minimization conditions, we get the expression for $\mu^{2}_{\Delta}$ as shown below,
$\mu^{2}_{\Delta}$ $=$ $\frac{1}{v_{2}}(-2\mu v_{H}^{2}-2\lambda^{H\Delta}_{3}v_{H}^{2}v_{1}-2\lambda_{4}^{H\Delta}v_{H}^{2}v_{1}-3\lambda_{1}^{H\Delta}v_{H}^{2}v_{2}+i(\sqrt{3}\lambda_{2}^{\Delta}v_{1}^{2}v_{2}$
$-\sqrt{3}\lambda_{2}^{\Delta}v_{2}^{3})-2\lambda_{1}^{\Delta}v_{1}^{2}v_{2}+\lambda_{2}^{\Delta}v_{1}^{2}v_{2}-4\lambda^{\Delta}_{3}v_{1}^{2}v_{2}-2\lambda_{4}^{\Delta}v_{1}^{2}v_{2}-2\lambda_{1}^{\Delta}v_{2}^{3}$
$+\lambda_{2}^{\Delta}v_{2}^{3}-\lambda_{1}^{\Delta\eta}v_{2}v_{\eta}^{2}-\lambda_{1}^{\Delta\kappa}v_{2}v_{\kappa}^{2}-\lambda_{1}^{\Delta\xi}v_{2}v_{\xi}^{2}-\lambda_{1}^{\Delta\chi}v_{2}v_{\chi}^{2}-\lambda_{1}^{\Delta\psi}v_{2}v_{\psi}).$
As $v_{1}$ and $v_{2}$ are real in our model, it is evident from the above expression that $\mu^{2}_{\Delta}$ is complex quantity. In order to restrict this, we have only one choice i.e., $\lambda_{2}^{\Delta}=0$. In addition, we obtain the following significant relations:
@@
$\left(\frac{v}{v_{2}}\right)^{2}$ $=$ $-\frac{2\left(\frac{v_{1}}{v_{2}}\right)^{3}\left(\lambda_{4}^{\Delta}+2\lambda_{3}^{\Delta}\left(\left(\frac{v_{1}}{v_{2}}\right)^{2}-1\right)+\lambda_{4}^{\Delta}\left(\frac{v_{1}}{v_{2}}\right)^{2}\right)}{3\,\lambda_{2}^{H\Delta}\left(1+\left(\frac{v_{1}}{v_{2}}\right)\left(\left(\frac{v_{1}}{v_{2}}\right)\left(\left(\frac{v_{1}}{v_{2}}\right)^{3}-2\left(\frac{v_{1}}{v_{2}}\right)+4\right)-3\right)\right)}.$ (9)
@@ 4.4 Refined Parameter Space
We first try to realize the parameter domain of the elements : $\lambda_{4}^{H\Delta}$, $\mu v_{2}$ and $(v/v_{2})^{2}$ which depend on the free parameters $\lambda_{3}^{\Delta}$, $\lambda_{4}^{\Delta}$, $\lambda_{2}^{H\Delta}$, $\lambda_{4}^{H\Delta}$ and $v_{1}/v_{2}$. The parameter $v_{1}/v_{2}$ is guided by $r$ and $\theta$. From the phenomenological analysis in section 2, we are aware of the parameter space of the two said texture parameters. Moreover, we adhere to the fact that the quartic coupling constants, $\lambda\sim\mathcal{O}(1)$. Motivated by this, we allow the free parameters $\lambda_{3}^{H\Delta}$, $\lambda_{3}^{\Delta}$, $\lambda_{4}^{\Delta}$ to vary in the range $(0.01-0.99)$. With these inputs, we graphically visualise the parameter space of $\lambda_{4}^{H\Delta}$, $\mu v_{2}$ and $(v/v_{2})^{2}$ in Figs. (6). We further extend the analysis to gain deeper insight of domain of $v_{1}$, $v_{2}$ and $\mu$ (see Fig. (7)).
@@ 5 Summary and Discussion
The study of neutrino mass matrix texture has been explored exhaustively by the model builders. At the same time, the experiments are gradually achieving more precise predictions, whereas the fundamental issues such as the neutrino mass hierarchy and anomalies related to physical parameters remain unresolved. Hence, there lies the importance of looking into more predictive textures of neutrino mass matrix both from a bottom-up and top-down perspective. Although physics is independent of parametrization, yet a meaningful parametrization may simplify a certain problem and thereby will help to approach a framework in a better way. With this motivation, the present piece of work adapts “Exponential parametrization”, where the general neutrino mass matrix element is visualised in polar form. We have posited a neutrino mass matrix texture under the roof of the said parametrisation which is quite promising in the light of neutrino oscillation experiments. We predict narrow bound/ allowed and forbidden regions for the observable parameters like atmospheric angle and three CP violating phases. The texture is perfectly obtained in its exact form ensuring the independence of the parameters, in a discrete symmetry based framework. Moreover in this work, we have emphasized on how accurately the texture is consistent with the model in terms of refined parameter space.
It is important to mention that a similar approach encompassing texture analysis and preference for certain parametrization may also assist us in exploring both quark and lepton sectors in the unified frameworks.
@@ Appendix A Product Rules of $A_{4}$
The discrete group $A_{4}$ is a subgroup of $SU(3)$ and has 12 elements. The group has four irreducible representation out of which three are one dimensional and one is three dimensional. In S-basis, the product of two triplets gives,
$3\times 3=1+1^{\prime}+1^{\prime\prime}+3_{S}+3_{A},$ (10)
@@
11 11 $=$ $(a_{1}\bar{b_{1}}+a_{2}\bar{b_{2}}+a_{3}\bar{b_{3}})$ (11)
$1^{\prime}$ $=$ $(a_{1}\bar{b_{1}}+\omega^{2}a_{2}\bar{b_{2}}+\omega a_{3}\bar{b_{3}})$ (12)
$1^{\prime\prime}$ $=$ $(a_{1}\bar{b_{1}}+\omega a_{2}\bar{b_{2}}+\omega^{2}a_{3}\bar{b_{3}})$ (13)
$(3\times 3)_{S}$ $=$ $\begin{bmatrix}a_{2}b_{3}+a_{3}b_{2}\\ a_{3}b_{1}+a_{1}b_{3}\\ a_{1}b_{2}+a_{2}b_{1}\\ \end{bmatrix},$ (14)
$(3\times 3)_{A}$ $=$ $\begin{bmatrix}a_{2}b_{3}-a_{3}b_{2}\\ a_{3}b_{1}-a_{1}b_{3}\\ a_{1}b_{2}-a_{2}b_{1}\\ \end{bmatrix}.$ (15)
@@ Appendix B The Scalar Potential
To justify the vacuum alignments of the scalar fields, we construct the $SU(2)_{L}\times U(1)_{Y}\times A_{4}\times Z_{10}$ invariant scalar potential in the following way,
$V$ $=$ $V(H)+V(\Delta)+V(\chi)+V(\psi)+V(\xi)+V(\eta)+V(\kappa)+V(H,\Delta)+V(H,\chi)+V(H,\psi)+V(H,\xi)$ (17)
$+V(H,\eta)+V(H,\kappa)+V(\Delta,\chi)+V(\Delta,\psi)+V(\Delta,\xi)+V(\Delta,\eta)+V(\Delta,\kappa)+V(\chi,\psi)+V(\chi,\xi)$
$+V(\chi,\eta)+V(\chi,\kappa)+V(\psi,\xi)+V(\psi,\eta)+V(\psi,\kappa)+V(\xi,\eta)+V(\xi,\kappa)+V(\eta,\kappa).$
@@
$V(H)$ $=$ $\mu^{2}_{H}(H^{\dagger}H)+\lambda^{H}_{1}(H^{\dagger}H)(H^{\dagger}H)+\lambda^{H}_{2}(H^{\dagger}H)_{1^{\prime}}(H^{\dagger}H)_{1^{\prime\prime}}+\lambda^{H}_{3}(H^{\dagger}H)_{3_{s}}(H^{\dagger}H)_{3_{s}}+\lambda^{H}_{4}(H^{\dagger}H)_{3_{s}}$ (18)
$(H^{\dagger}H)_{3_{a}}+\lambda^{H}_{5}(H^{\dagger}H)_{3_{a}}(H^{\dagger}H)_{3_{a}},$
$V(\Delta)$ $=$ $\mu^{2}_{\Delta}Tr(\Delta^{\dagger}\Delta)+\lambda^{\Delta}_{1}Tr(\Delta^{\dagger}\Delta)Tr(\Delta^{\dagger}\Delta)+\lambda^{\Delta}_{2}Tr(\Delta^{\dagger}\Delta)_{1^{\prime}}Tr(\Delta^{\dagger}\Delta)_{1^{\prime\prime}}+\lambda^{\Delta}_{3}Tr(\Delta^{\dagger}\Delta)_{3_{s}}Tr(\Delta^{\dagger}\Delta)_{3_{s}}$ (19)
$+\lambda^{\Delta}_{4}Tr(\Delta^{\dagger}\Delta)_{3_{s}}Tr(\Delta^{\dagger}\Delta)_{3_{a}}+\lambda^{\Delta}_{5}Tr(\Delta^{\dagger}\Delta)_{3_{a}}Tr(\Delta^{\dagger}\Delta)_{3_{a}},$
$V(\chi)$ $=$ $\mu^{2}_{\chi}(\chi^{\dagger}\chi)+\lambda^{\chi}(\chi^{\dagger}\chi)(\chi^{\dagger}\chi),$ (20)
$V(\psi)$ $=$ $\mu^{2}_{\psi}(\psi^{\dagger}\psi)+\lambda^{\psi}(\psi^{\dagger}\psi)(\psi^{\dagger}\psi),$ (21)
$V(\xi)$ $=$ $\mu^{2}_{\xi}(\xi^{\dagger}\xi)+\lambda^{\xi}(\xi^{\dagger}\xi)(\xi^{\dagger}\xi),$ (22)
$V(\eta)$ $=$ $\mu^{2}_{\eta}(\eta^{\dagger}\eta)+\lambda^{\eta}(\eta^{\dagger}\eta)(\eta^{\dagger}\eta),$ (23)
$V(\kappa)$ $=$ $\mu^{2}_{\kappa}(\kappa^{\dagger}\kappa)+\lambda^{\kappa}(\kappa^{\dagger}\kappa)(\kappa^{\dagger}\kappa),$ (24)
$V(H,\Delta)$ $=$ $\lambda^{H\Delta}_{1}(H^{\dagger}H)Tr(\Delta^{\dagger}\Delta)+\lambda^{H\Delta}_{2}[(H^{\dagger}H)_{1^{\prime}}Tr(\Delta^{\dagger}\Delta)_{1^{\prime\prime}}+(H^{\dagger}H)_{1^{\prime\prime}}Tr(\Delta^{\dagger}\Delta)_{1^{\prime}}]+\lambda^{H\Delta}_{3}(H^{\dagger}H)_{3_{s}}$ (25)
$Tr(\Delta^{\dagger}\Delta)_{3_{s}}+\lambda^{H\Delta}_{4}[(H^{\dagger}H)_{3_{s}}Tr(\Delta^{\dagger}\Delta)_{3_{a}}+(H^{\dagger}H)_{3_{a}}Tr(\Delta^{\dagger}\Delta)_{3_{s}}]+\lambda^{H\Delta}_{5}(H^{\dagger}H)_{3_{a}}$
$Tr(\Delta^{\dagger}\Delta)_{3_{a}}+\mu(H^{T}i\sigma_{2}\Delta^{\dagger}H),$
$V(H,\chi)$ $=$ $\lambda^{H\chi}_{1}(H^{\dagger}H)(\chi^{\dagger}\chi),$ (26)
$V(H,\psi)$ $=$ $\lambda^{H\psi}_{1}(H^{\dagger}H)(\psi^{\dagger}\psi),$ (27)
$V(H,\xi)$ $=$ $\lambda^{H\xi}_{1}(H^{\dagger}H)(\xi^{\dagger}\xi),$ (28)
$V(H,\eta)$ $=$ $\lambda^{H\eta}_{1}(H^{\dagger}H)(\eta^{\dagger}\eta),$ (29)
$V(H,\kappa)$ $=$ $\lambda^{H\kappa}_{1}(H^{\dagger}H)(\kappa^{\dagger}\kappa),$ (30)
$V(\Delta,\chi)$ $=$ $\lambda^{\Delta\chi}_{1}Tr(\Delta^{\dagger}\Delta)(\chi^{\dagger}\chi),$ (31)
$V(\Delta,\psi)$ $=$ $\lambda^{\Delta\psi}_{1}Tr(\Delta^{\dagger}\Delta)(\psi^{\dagger}\psi),$ (32)
$V(\Delta,\xi)$ $=$ $\lambda^{\Delta\xi}_{1}Tr(\Delta^{\dagger}\Delta)(\xi^{\dagger}\xi),$ (33)
$V(\Delta,\eta)$ $=$ $\lambda^{\Delta\eta}_{1}Tr(\Delta^{\dagger}\Delta)(\eta^{\dagger}\eta),$ (34)
$V(\Delta,\kappa)$ $=$ $\lambda^{\Delta\kappa}_{1}Tr(\Delta^{\dagger}\Delta)(\kappa^{\dagger}\kappa),$ (35)
$V(\chi,\psi)$ $=$ $\lambda^{\chi\psi}_{1}(\chi^{\dagger}\chi)(\psi^{\dagger}\psi),$ (36)
$V(\chi,\xi)$ $=$ $\lambda^{\chi\xi}_{1}(\chi^{\dagger}\chi)(\xi^{\dagger}\xi),$ (37)
$V(\chi,\eta)$ $=$ $\lambda^{\chi\eta}_{1}(\chi^{\dagger}\chi)(\eta^{\dagger}\eta),$ (38)
$V(\chi,\kappa)$ $=$ $\lambda^{\chi\kappa}_{1}(\chi^{\dagger}\chi)(\kappa^{\dagger}\kappa),$ (39)
$V(\psi,\xi)$ $=$ $\lambda^{\psi\xi}_{1}(\psi^{\dagger}\psi)(\xi^{\dagger}\xi),$ (40)
$V(\psi,\eta)$ $=$ $\lambda^{\psi\eta}_{1}(\psi^{\dagger}\psi)(\eta^{\dagger}\eta),$ (41)
$V(\psi,\kappa)$ $=$ $\lambda^{\psi\kappa}_{1}(\psi^{\dagger}\psi)(\kappa^{\dagger}\kappa),$ (42)
$V(\xi,\eta)$ $=$ $\lambda^{\xi\eta}_{1}(\xi^{\dagger}\xi)(\eta^{\dagger}\eta),$ (43)
$V(\xi,\kappa)$ $=$ $\lambda^{\xi\kappa}_{1}(\xi^{\dagger}\xi)(\kappa^{\dagger}\kappa),$ (44)
$V(\eta,\kappa)$ $=$ $\lambda^{\eta\kappa}_{1}(\eta^{\dagger}\eta)(\kappa^{\dagger}\kappa).$ (45)

### v3

@@ 4.2 Important Relations
$\frac{v_{\chi}^{2}}{v_{\eta}}$ $=$ $\frac{\Lambda^{2}y_{c}\left(b_{1}+\sqrt{3}b_{2}+g\right)^{2}}{6v^{2}y_{2}^{2}(-2b_{1}+2ib_{2}+f)},$ (5)
$v_{\xi}$ $=$ $-\frac{6\sqrt{3}v^{2}y_{1}^{2}}{\sqrt{3}ay_{a}-2\sqrt{3}b_{1}y_{a}-6b_{2}y_{a}},$ (6)
where, $b_{1}=r\cos\theta$ and $b_{2}=r\sin\theta$. We expect that these relations, somehow, will help to understand the numerical domain of the model parameters more closely. The parameter space of the elements, $v_{1}/v_{2}$, $|v_{\chi}^{2}/v_{\eta}|$ and $|v_{\xi}|$ are presented in Fig. 5). For this, we allow the cut-off scale and the Yukawa couplings to vary randomly in the bounds: $(10^{3}-10^{16})$ GeV and (0.010.99)(0.01-0.99)( 0.01 - 0.99 ) respectively. Needless to mention, the $v_{H}$ is taken as 246246246 GeV.
Figure 5: Shows the correlation plots between (a) $v_{1}/v_{2}$ vs $|v_{\xi}|$. (b) $\log(|v_{\xi}|)$ vs $\log(|v_{\chi}^{2}/v_{\eta}|)$. (c) $\log(\Lambda)$ vs $\log(|v_{\xi}|)$. The red points signify texture prediction (TP) while the blue data points signify the corrections from model sector(MP).
@@ 4.3 The Scalar Sector
There are seven $A_{4}$ complex scalar field multiplets in the model and, the associated $SU(2)_{L}\times U(1)_{Y}\times A_{4}\times Z_{10}\times Z_{7}$ invariant potential, is shown explicitly in B. We see that there are seven $\mu^{2}$’s, forty $\lambda$’s and one $\mu$ in the potential expression. Thus it enhances the liberty to choose the desired vacuum alignments for the scalar fields. The eleven minimisation conditions in support of the vacuum alignments as discussed in section 4.1 are elaborated in B.
From the minimization conditions, we obtain the following significant relations:
@@
$\left(\frac{v}{v_{2}}\right)^{2}$ $=$ $\frac{\left(\frac{v_{1}}{v_{2}}\right)^{3}\left(\left(\left(\frac{v_{1}}{v_{2}}\right)^{2}-1\right)(3\lambda_{2}^{\Delta}-4\lambda_{3}^{\Delta})-2\lambda_{4}^{\Delta}\left(\left(\frac{v_{1}}{v_{2}}\right)^{2}+1\right)\right)}{3\left(\lambda_{2}^{H\Delta}+\lambda_{2}^{H\Delta}\left(\frac{v_{1}}{v_{2}}\right)\left(\left(\frac{v_{1}}{v_{2}}\right)\left(\left(\frac{v_{1}}{v_{2}}\right)^{3}-2\left(\frac{v_{1}}{v_{2}}\right)+4\right)-3\right)\right)}.$ (9)
@@ 4.4 Refined Parameter Space
We first try to realize the parameter domain of the elements : $\lambda_{4}^{H\Delta}$, $\mu v_{2}$ and $(v/v_{2})^{2}$ which depend on the free parameters $\lambda_{2}^{\Delta}$, $\lambda_{3}^{\Delta}$, $\lambda_{4}^{\Delta}$, $\lambda_{2}^{H\Delta}$, $\lambda_{4}^{H\Delta}$ and $v_{1}/v_{2}$. The parameter $v_{1}/v_{2}$ is guided by $r$ and $\theta$. From the phenomenological analysis in section 2, we are aware of the parameter space of the two said texture parameters. Moreover, we adhere to the fact that the quartic coupling constants, $\lambda\sim\mathcal{O}(1)$. Motivated by this, we allow the free parameters $\lambda_{3}^{H\Delta}$, $\lambda_{2}^{\Delta}$, $\lambda_{3}^{\Delta}$, $\lambda_{4}^{\Delta}$ to vary in the range (0.010.99)(0.01-0.99)( 0.01 - 0.99 ). With these inputs, we graphically visualise the parameter space of $\lambda_{4}^{H\Delta}$, $\mu v_{2}$ and $(v/v_{2})^{2}$ in Figs. (6). We further extend the analysis to gain deeper insight of domain of $v_{1}$, $v_{2}$ and $\mu$ (see Fig. (7)).
@@ 5 Summary and Discussion
The study of neutrino mass matrix texture has been explored exhaustively by the model builders. At the same time, the experiments are gradually achieving more precise predictions, whereas the fundamental issues such as the neutrino mass hierarchy and anomalies related to physical parameters remain unresolved. Hence, there lies the importance of looking into more predictive textures of neutrino mass matrix both from a bottom-up and top-down perspective. Although physics is independent of parametrization, yet a meaningful parametrization may simplify a certain problem and thereby will help to approach a framework in a better way. With this motivation, the present piece of work adapts “Exponential parametrization”, where the general neutrino mass matrix element is visualised in polar form. We have posited a neutrino mass matrix texture under the roof of the said parametrisation which is quite promising in the light of neutrino oscillation experiments. We predict narrow bound/ allowed and forbidden regions for the observable parameters like atmospheric angle and three CP violating phases. The texture is perfectly obtained in its exact form ensuring the independence of the parameters, in a discrete symmetry based framework. Moreover in this work, we have emphasized on how accurately the texture is consistent with the model in terms of refined parameter space. It is important to mention that a similar approach encompassing texture analysis and preference for certain parametrization may also assist us in exploring both quark and lepton sectors in the unified frameworks.
In the context of scalar potential and minimization conditions, we would like to state that the strategy is not so straightforward. The present work deals with seven scalar field multiplets as shown in Table (2), and in reality, there are twenty-nine real scalar fields. Hence, we obtain twenty-nine optimisation conditions. To understand whether the chosen vacuum alignment leads to the actual minima, the Hessian matrix: $H_{ij}=(\frac{\partial^{2}V}{\partial\phi_{i}\partial\phi_{j}})_{0}$, is to be evaluated. The necessary condition for a minimum to occur is to ensure that either the eigenvalues of $H$ are positive (or zero for Goldstone states), or the principal cofactors are positive definite. In the present context, this task is quite challenging, as the Hessian matrix is of dimension twenty-nine. However, we have considered the optimization conditions as the minimization conditions based on the fact that the potential (or the Hessian) involves forty-eight free parameters: one $\mu$, seven $\mu^{2}$’s, and forty $\lambda$’s. By tuning these parameters appropriately, one may achieve the required conditions. However, this task is beyond the scope of the present work.
@@ Appendix A Product Rules of $A_{4}$
The discrete group $A_{4}$ is a subgroup of $SU(3)$ and has 12 elements. The group has four irreducible representation out of which three are one dimensional and one is three dimensional. In S-basis, the product of two triplets $(a_{1},a_{2},a_{3})^{T}$ and $(b_{1},b_{2},b_{3})^{T}$ gives,
$3\times 3=1+1^{\prime}+1^{\prime\prime}+3_{S}+3_{A},$
@@
$(3\times 3)_{1}$ $=$ $(a_{1}b_{1}+a_{2}b_{2}+a_{3}b_{3}),$
$(3\times 3)_{1^{\prime}}$ $=$ $(a_{1}b_{1}+\omega^{2}a_{2}b_{2}+\omega a_{3}b_{3}),$
$(3\times 3)_{1^{\prime\prime}}$ $=$ $(a_{1}b_{1}+\omega a_{2}b_{2}+\omega^{2}a_{3}b_{3}),$
$(3\times 3)_{3_{S}}$ $=$ $\begin{bmatrix}a_{2}b_{3}+a_{3}b_{2}\\ a_{3}b_{1}+a_{1}b_{3}\\ a_{1}b_{2}+a_{2}b_{1}\\ \end{bmatrix},$
$(3\times 3)_{3_{A}}$ $=$ $\begin{bmatrix}a_{2}b_{3}-a_{3}b_{2}\\ a_{3}b_{1}-a_{1}b_{3}\\ a_{1}b_{2}-a_{2}b_{1}\\ \end{bmatrix}.$
@@ Appendix B The Scalar Potential
To justify the vacuum alignments of the scalar fields, we construct the $SU(2)_{L}\times U(1)_{Y}\times A_{4}\times Z_{10}\times Z_{7}$ invariant scalar potential in the following way,
$V$ $=$ $V(H)+V(\Delta)+V(\chi)+V(\psi)+V(\xi)+V(\eta)+V(\kappa)+V(H,\Delta)$
$+V(H,\chi)+V(H,\psi)+V(H,\xi)+V(H,\eta)+V(H,\kappa)+V(\Delta,\chi)$
$+V(\Delta,\psi)+V(\Delta,\xi)+V(\Delta,\eta)+V(\Delta,\kappa)+V(\chi,\psi)+V(\chi,\xi)$
$+V(\chi,\eta)+V(\chi,\kappa)+V(\psi,\xi)+V(\psi,\eta)+V(\psi,\kappa)+V(\xi,\eta)$
$+V(\xi,\kappa)+V(\eta,\kappa)+\,h.c.$
@@
$V(H)$ $=$ $\mu^{2}_{H}(H^{\dagger}H)+\lambda^{H}_{1}(H^{\dagger}H)(H^{\dagger}H)+\lambda^{H}_{2}(H^{\dagger}H)_{1^{\prime}}(H^{\dagger}H)_{1^{\prime\prime}}+$
$\lambda^{H}_{3}(H^{\dagger}H)_{3_{s}}(H^{\dagger}H)_{3_{s}}+\lambda^{H}_{4}(H^{\dagger}H)_{3_{s}}(H^{\dagger}H)_{3_{a}}+\lambda^{H}_{5}(H^{\dagger}$
$H)_{3_{a}}(H^{\dagger}H)_{3_{a}},$
$V(\Delta)$ $=$ $\mu^{2}_{\Delta}Tr(\Delta^{\dagger}\Delta)+\lambda^{\Delta}_{1}Tr(\Delta^{\dagger}\Delta)Tr(\Delta^{\dagger}\Delta)+\lambda^{\Delta}_{2}Tr(\Delta^{\dagger}\Delta)_{1^{\prime}}$
$Tr(\Delta^{\dagger}\Delta)_{1^{\prime\prime}}+\lambda^{\Delta}_{3}Tr(\Delta^{\dagger}\Delta)_{3_{s}}Tr(\Delta^{\dagger}\Delta)_{3_{s}}+\lambda^{\Delta}_{4}Tr(\Delta^{\dagger}\Delta)_{3_{s}}$
$Tr(\Delta^{\dagger}\Delta)_{3_{a}}+\lambda^{\Delta}_{5}Tr(\Delta^{\dagger}\Delta)_{3_{a}}Tr(\Delta^{\dagger}\Delta)_{3_{a}},$
$V(\chi)$ $=$ $\mu^{2}_{\chi}(\chi^{\dagger}\chi)+\lambda^{\chi}(\chi^{\dagger}\chi)(\chi^{\dagger}\chi),$
$V(\psi)$ $=$ $\mu^{2}_{\psi}(\psi^{\dagger}\psi)+\lambda^{\psi}(\psi^{\dagger}\psi)(\psi^{\dagger}\psi),$
$V(\xi)$ $=$ $\mu^{2}_{\xi}(\xi^{\dagger}\xi)+\lambda^{\xi}(\xi^{\dagger}\xi)(\xi^{\dagger}\xi),$
$V(\eta)$ $=$ $\mu^{2}_{\eta}(\eta^{\dagger}\eta)+\lambda^{\eta}(\eta^{\dagger}\eta)(\eta^{\dagger}\eta),$
$V(\kappa)$ $=$ $\mu^{2}_{\kappa}(\kappa^{\dagger}\kappa)+\lambda^{\kappa}(\kappa^{\dagger}\kappa)(\kappa^{\dagger}\kappa),$
$V(H,\Delta)$ $=$ $\lambda^{H\Delta}_{1}(H^{\dagger}H)Tr(\Delta^{\dagger}\Delta)+\lambda^{H\Delta}_{2}[(H^{\dagger}H)_{1^{\prime}}Tr(\Delta^{\dagger}\Delta)_{1^{\prime\prime}}+$
$(H^{\dagger}H)_{1^{\prime\prime}}Tr(\Delta^{\dagger}\Delta)_{1^{\prime}}]+\lambda^{H\Delta}_{3}(H^{\dagger}H)_{3_{s}}Tr(\Delta^{\dagger}\Delta)_{3_{s}}+\lambda^{H\Delta}_{4}$
$[(H^{\dagger}H)_{3_{s}}Tr(\Delta^{\dagger}\Delta)_{3_{a}}+(H^{\dagger}H)_{3_{a}}Tr(\Delta^{\dagger}\Delta)_{3_{s}}]+\lambda^{H\Delta}_{5}(H^{\dagger}$
$H)_{3_{a}}Tr(\Delta^{\dagger}\Delta)_{3_{a}}+\frac{\mu_{H\Delta}}{\Lambda}(H^{T}i\sigma_{2}\Delta^{T}H\xi),$
$V(H,\chi)$ $=$ $\lambda^{H\chi}_{1}(H^{\dagger}H)(\chi^{\dagger}\chi),$
$V(H,\psi)$ $=$ $\lambda^{H\psi}_{1}(H^{\dagger}H)(\psi^{\dagger}\psi),$
$V(H,\xi)$ $=$ $\lambda^{H\xi}_{1}(H^{\dagger}H)(\xi^{\dagger}\xi),$
$V(H,\eta)$ $=$ $\lambda^{H\eta}_{1}(H^{\dagger}H)(\eta^{\dagger}\eta),$
$V(H,\kappa)$ $=$ $\lambda^{H\kappa}_{1}(H^{\dagger}H)(\kappa^{\dagger}\kappa),$
$V(\Delta,\chi)$ $=$ $\lambda^{\Delta\chi}_{1}Tr(\Delta^{\dagger}\Delta)(\chi^{\dagger}\chi),$
$V(\Delta,\psi)$ $=$ $\lambda^{\Delta\psi}_{1}Tr(\Delta^{\dagger}\Delta)(\psi^{\dagger}\psi),$
$V(\Delta,\xi)$ $=$ $\lambda^{\Delta\xi}_{1}Tr(\Delta^{\dagger}\Delta)(\xi^{\dagger}\xi),$
$V(\Delta,\eta)$ $=$ $\lambda^{\Delta\eta}_{1}Tr(\Delta^{\dagger}\Delta)(\eta^{\dagger}\eta),$
$V(\Delta,\kappa)$ $=$ $\lambda^{\Delta\kappa}_{1}Tr(\Delta^{\dagger}\Delta)(\kappa^{\dagger}\kappa),$
$V(\chi,\psi)$ $=$ $\lambda^{\chi\psi}_{1}(\chi^{\dagger}\chi)(\psi^{\dagger}\psi),$
$V(\chi,\xi)$ $=$ $\lambda^{\chi\xi}_{1}(\chi^{\dagger}\chi)(\xi^{\dagger}\xi),$
$V(\chi,\eta)$ $=$ $\lambda^{\chi\eta}_{1}(\chi^{\dagger}\chi)(\eta^{\dagger}\eta),$
$V(\chi,\kappa)$ $=$ $\lambda^{\chi\kappa}_{1}(\chi^{\dagger}\chi)(\kappa^{\dagger}\kappa),$
$V(\psi,\xi)$ $=$ $\lambda^{\psi\xi}_{1}(\psi^{\dagger}\psi)(\xi^{\dagger}\xi),$
$V(\psi,\eta)$ $=$ $\lambda^{\psi\eta}_{1}(\psi^{\dagger}\psi)(\eta^{\dagger}\eta),$
$V(\psi,\kappa)$ $=$ $\lambda^{\psi\kappa}_{1}(\psi^{\dagger}\psi)(\kappa^{\dagger}\kappa),$
$V(\xi,\eta)$ $=$ $\lambda^{\xi\eta}_{1}(\xi^{\dagger}\xi)(\eta^{\dagger}\eta),$
$V(\xi,\kappa)$ $=$ $\lambda^{\xi\kappa}_{1}(\xi^{\dagger}\xi)(\kappa^{\dagger}\kappa),$
$V(\eta,\kappa)$ $=$ $\lambda^{\eta\kappa}_{1}(\eta^{\dagger}\eta)(\kappa^{\dagger}\kappa).$